OpenStudy (shamim):

integrate dx/(1-cosx)

OpenStudy (jhannybean):

dammit!!!! omg you read my mind lmao.

hartnn (hartnn):

another way is to use double angle formula. cos 2x =... ?

hartnn (hartnn):

@shamim several methods are being shown here, which way you would like to proceed ?

OpenStudy (primeralph):

I dropped mine. Only gets more convoluted.

OpenStudy (shamim):

confused

hartnn (hartnn):

do u know integration of cosec^2 x ?

OpenStudy (shamim):

no

hartnn (hartnn):

i'll tell you before that can you tell me all formulas for cos 2x that you know ?

OpenStudy (jhannybean):

\[\large \int \cfrac{1}{1- \cos(x)}dx\]\[\large \int \cfrac{1}{1- \cos(x)} \cdot \cfrac{1+ \cos(x)}{1+ \cos(x)}\]\[\large \int \cfrac{1+ \cos(x)}{1- \cos^2 (x)}dx\]\[\large \int \cfrac{1+ \cos(x)}{ \sin^2 (x)}dx\]

OpenStudy (shamim):

go on @Jhannybean

OpenStudy (shamim):

i m getting ur solution

OpenStudy (shamim):

@Jhannybean

OpenStudy (jhannybean):

rewrite \( \sin^2(x)\) as \( \csc^2 (x)\) and integrate.

OpenStudy (shamim):

please show every steps

OpenStudy (shamim):

sorry

OpenStudy (jhannybean):

Try it out. Integrate \[\large \int (1+ \cos(x))( \csc^2 (x))dx\] You can do it :) Tell me what you get and i'll check with my methods of solving the problem

OpenStudy (shamim):

then

OpenStudy (shamim):

i m not able to do this integration

OpenStudy (jhannybean):

Yes,you can do it... distribute csc^2(x).What do you get?

OpenStudy (shamim):

bt i understood ur steps

OpenStudy (shamim):

show plz

OpenStudy (jhannybean):

I've got to head off the computer, @hartnn can help you finish :D

hartnn (hartnn):

@shamim why don't you atleast give a shot, whats (cosec^2 x)(1+cos x) =... ? u know basic trigonometric identities, right ?

OpenStudy (shamim):

\[cosec ^{2}x+cosec ^{2}x cosx\]

OpenStudy (anonymous):

From WolframAlpha: